# The Dynamic, the Static, and the Weak factor models and the analysis of high-dimensional time series

### v1

@@ 5 The Weak: De Mol, Giannone, and Reichlin, Onatski, Hallin and Liška, and Gersing, Rust, and Deistler
    Statically rate-weak factors in the approximate static factor model are introduced by De Mol et al. (2008) and Onatski (2012). They are related with the eigenvalues $\lambda^{(n)}_{\chi;j}$, $j=1,\ldots,r$, of the common component covariance matrix $\bm{\Gamma}_{\chi}$ which are diverging at sublinear rates, usually parametrized as $n^{\alpha_{j}}$ with $\alpha_{j}<1$. This is the most popular setting, see Lam and Yao (2012), Freyaldenhoven (2022), Uematsu and Yamagata (2022), and Bai and Ng (2023), among many others.
@@
    Dynamically rate-weak factors in the GDFM are considered by Barigozzi and Farnè (2024) who proposed an estimator of a low-rank plus sparse spectral density matrix, under the assumption that the low-rank component has eigenvalues $\lambda^{(n)}_{\chi;j}(\theta)$, $j=1,\ldots,q$, diverging at sublinear rates $n^{\alpha_{j}}$ with $\alpha_{j}<1$, $\theta\text{-a.e. in }(-\pi,\pi]$.
@@
    Block-specific factors are introduced by Hallin and Liška (2011), under a GDFM approach, in the context of panels divided into subpanels or blocks. The terminology weakly common factors is used for dynamic factors that are dynamically common in some block(s) but dynamically idiosyncratic in some other(s). The associated spectral eigenvalues are assumed to be diverging with the dimension of the block but no rate is specified. Approximate static factor models with a hierarchical block structure are considered, for example, by Moench et al. (2013) but without any explicit mention of weak factors.
@@
    Finally, Gersing et al. (2023) define weakly common components as the difference (at time $t$) between the dynamically common and the statically common components of $X_{it}$. Any (possibly infinite-dimensional) orthonormal basis of the space they are spanning (at time $t$) then can be considered a basis of weak factors at time $t$ call them GRD-weak factors. This notion of weakness bears no relation to (a) and (b) above.
@@ 5.1 Rate-weak factors: De Mol, Giannone and Reichlin, Onatski
As mentioned in the previous sections, as long as we have a widening, as $n\to\infty$, of the eigen-gap between the first $q$ or $r$ eigenvalues of the spectral density or covariance matrix of the observables $\mathbf{X}^{(n)}$ and the remaining $n-q$ or $n-r$ ones, the GDFM or the approximate static factor models are identified. Thus, for identification purposes, there is no need to assume any specific divergence rate for the divergent eigenvalues.
@@ 5.2 Weak factors in panels with block structure: Hallin and Liška
A similar approach could be taken, mutatis mutandis, with similar comments, in the context of approximate static factor models, yielding a refinement of the static factor model decomposition (2.1).
@@ 5.3 Weak factors are everywhere: Gersing, Rust, and Deistler
Under the intriguing title “Weak factors are everywhere,” Gersing et al. (2023), with the objective of reconciling the Static and the Dynamic factors, show (their approach slightly differs from the one developed here) that the statically common space $\mathcal{H}_{\scriptscriptstyle\text{\rm stat com}}^{{\bf X}_{t}}$, for any $t$, is a subspace of the dynamically common one $\mathcal{H}_{\scriptscriptstyle\text{\rm dyn com}}^{{\bf X}}$ and call weakly common at time $t$ the elements of the difference $\mathcal{H}_{\scriptscriptstyle\text{\rm dyn com}}^{{\bf X}}\setminus\mathcal{H}_{\scriptscriptstyle\text{\rm stat com}}^{{\bf X}_{t}}$: we shall call them GRD-weak.
This result naturally yields a new canonical factor decomposition, of the form
@@
where $\chi_{it}^{\scriptscriptstyle\text{\rm weak}}$—the GRD-weakly common component of $\mathbf{X}_{it}$—is the difference between the dynamically and statically common components $\chi_{it}^{\scriptscriptstyle\text{\rm dyn}}$ and $\chi_{it}^{\scriptscriptstyle\text{\rm stat}}$ of $X_{it}$. This is a very ingenious idea, which shows how the GDFM (as defined in (3.5)) is overarching the static model (as defined in (2.1)) and, with the additional assumption (4.3), strictly nests the popular static model of Chamberlain (1983) and Chamberlain and Rothschild (1983). In both cases, the dynamically common space $\mathcal{H}_{\scriptscriptstyle\text{\rm dyn com}}^{{\bf X}}$ incorporates the GRD-weakly common components which, in the static approach are idiosyncratic, which explains the empirical finding (Forni et al., 2018) that the GDFM performs better in terms, e.g., of forecasting, than the static model (2.1) of Chamberlain and Rothschild even when the assumptions of the latter are satisfied.
@@
In this new canonical decomposition (5.1), being in the dynamically common space $\mathcal{H}_{\scriptscriptstyle\text{\rm dyn com}}^{{\bf X}}$, $\chi_{it}^{\scriptscriptstyle\text{\rm weak}}$ is orthogonal, at all leads and lags, to the dynamically idiosyncratic component ${\xi_{it}^{\scriptscriptstyle\text{\rm dyn}}}$; being in the statically idiosyncratic space at time $t$, it is also orthogonal to the statically common component at time $t$ ${\chi_{it}^{\scriptscriptstyle\text{\rm stat}}}$ (but not its leads and lags); that ${\chi_{it}^{\scriptscriptstyle\text{\rm stat}}}$ might even be a.s. zero (yielding a statically purely idiosyncratic process). Therefore, it cannot be a static factor, neither rate-strong nor rate-weak, and consists of non-pervasive lagged values of the (rate-weak or rate-strong) dynamic factors (lag-zero values included), or combinations thereof. This concept of GRD-weakly common component, thus is unrelated to the concepts of rate-weak factors discussed above.
@@ 6 Exchangeability: weak factors are nowhere!
When no block structure is present in the data, one may wonder whether the concept of rate-weak factors is really relevant in practice. An all too often overlooked feature of panel data, indeed, is that, unlike time-ordering, their cross-sectional ordering is entirely arbitrary. Very often, some alphabetical ordering is adopted, a choice that should not have any impact on the analysis. An observed panel, actually, should be treated as the equivalence class of its $n!$ cross-sectional permutations. In this section we consider for simplicity of notation only the static case. The same arguments can be made verbatim for the dynamic case at the price of heavier notation.
@@ 6.1 Factors and cross-sectional permutations: intuition
The left-hand plot in Figure 1 shows the evolution, for increasing $n=1,\ldots,n_{0}$, of the first eigenvalue $\lambda_{X;1}^{(n)}$ (normalized by $\lambda_{X;1}^{(n_{0})}$) of the $(n\times n)$ lag-zero covariance matrix ${\bm{\Gamma}}^{(n)}_{X}$ in the case of three cross-sectional permutations of the $X_{it}$’s. Let $B_{(i)}$, $i=1,\ldots,n_{0}$ denote the $i$th order statistic of the loadings: $B_{(1)}\leq B_{(2)}\leq\leavevmode\nobreak\ \!\ldots\leavevmode\nobreak\ \!\leq\leavevmode\nobreak\ \!B_{(n_{0})}$. The first permutation (dashed line) is ordering the cross-sectional items by increasing values of the loadings ($B_{i}=B_{(i)}$); the second permutation (dashed-dotted line) orders them by decreasing values ($B_{i}=B_{(n_{0}-i+1)}$); the third one (solid line) is alternating small and large values of the same ($B_{1}=B_{(n_{0})},B_{2}=B_{(1)},B_{3}=B_{(n_{0}-1)},\ldots,B_{n_{0}}=B_{(n_{0}/2)}$ for even $n_{0}$). From a naïve point of view, one would be tempted by a superlinear extrapolation in the first case (hence, a rate-superstrong factor) and a sublinear one (a rate-weak factor) in the second case, whereas a linear extrapolation (a rate-strong factor) looks quite reasonable in the third case. The three cases, however, are dealing with the same panel!
@@
Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption
Figure 3: Real data (McCracken and Ng, 2016), $r=8$. Evolution for increasing $n=1,\ldots,n_{0}$ of the eigenvalues $\lambda_{X;k}^{(n)}$, $k=1,\ldots,8$ (all normalized by $\lambda_{X;1}^{(n_{0})}$) under 100 random cross-sectional permutations.
Finally, we consider the dataset FRED-MD described by McCracken and Ng (2016) consisting of $n_{0}=126$ times series covering both the real and nominal sectors of the U.S. economy and including also labor market indicators, and financial variables. [^1] See https://research.stlouisfed.org/econ/mccracken/fred-databases/ The data is transformed to stationarity and missing values are imputed by means of the routines made available by McCracken and Ng (2016), which produces a balanced panel, with a sample running from April 1959 through March 2024, for an observation period of $T=780$ time points. The cross-sectional ordering, in this dataset, is fully arbitrary, hence meaningless. Standard methods (Bai and Ng, 2002) for determining the number of factors point to as much as $r=8$ factors. In Figure 3 we show the evolution for increasing $n=1,\ldots,n_{0}$ of the eigenvalues $\lambda_{X;k}^{(n)}$, $k=1,\ldots,8$ (all normalized by $\lambda_{X;1}^{(n_{0})}$) of ${\bm{\Gamma}}^{(n)}_{X}$ when considering 100 random permutations of the cross-section. Once again, we see that the evolution of all eigenvalues is always well described by a linear divergence in $n$, despite having different slopes due to the fact that eigenvalues are distinct.
Inspection of Figure 3 suggests that the last two or three factors are “weakly influential” (sufficiently so, however, for being detected by the Bai and Ng, 2002, method), and one may be tempted to consider them as rate-weak. Deciding about linear or sublinear divergence—in particular, testing the hypothesis of a rate-weak factor against rate-strong alternatives or estimating the value $\alpha$ in a sublinear rate $n^{\alpha}$—is extremely delicate. In view of Sections 6.2 and 7 below, the issue, however, is probably of limited practical importance.
@@ 6.2 Factors and cross-sectional permutations: exchangeability
Distributional symmetries and group invariance properties are among the most convincing, irrefutable, and forceful arguments giving statistical inference problems a structure which, when it exists, dramatically impacts its solutions. Invariance arguments have a long history in mathematical statistics (see Chapter 6 in Lehmann and Romano (2024)); their pertinence, in a broad variety of problems, has been highlighted recently in Austern and Orbanz (2022).
Stationarity is one of those fundamental invariance properties and, under our fully nonparametric approach (no parameters), is the only assumption we are imposing in order to obtain a factor model decomposition. Namely, we assume that the observed panel is the finite $(n\times T)$ realization of a stochastic process $\bf X$ of the form $\big{\{}X_{it}|\,i\in{\mathbb{N}},t\in{\mathbb{Z}}\big{\}}$ in the class $\mathcal{C}$ of real-valued second-order time-stationary processes indexed on ${\mathbb{N}}\times{\mathbb{Z}}$. Second-order stationarity imposes a restriction (stationarity) on time dependence, namely,
@@
The stochastic process $\big{\{}X_{it}|i\in{\mathbb{N}},t\in{\mathbb{Z}}\big{\}}$ is (second-order) cross-sectionally exchangeable—namely, for any $k\in{\mathbb{N}}$, any $k$-tuple $i_{1},\ldots,i_{k}$, and any permutation $\pi$ in the set $\Pi_{k}$ of the permutations of the integers $(1,\ldots,k)$, the distribution (the second-order moments) of the $k$-dimensional stochastic processes
$\big{\{}(X_{i_{1}t},\ldots,X_{i_{k}t}):\,t\in{\mathbb{Z}}\big{\}}\quad\text{and}\quad\big{\{}(X_{i_{{\pi(1)}}t},\ldots,X_{i_{{\pi(k)}}t}):\,t\in{\mathbb{Z}}\big{\}}$
@@
Clearly, exchangeability implies second-order exchangeability, viz. ${\rm E}\left[X_{i_{\ell}t}X_{i_{m}t^{\prime}}\right]={\rm E}\left[X_{i_{\pi(\ell)}t}X_{i_{\pi(m)}t^{\prime}}\right]$ for any $\ell$ and $m$ in $(1,\ldots,k)$, any $t$ and $t^{\prime}$ in $\mathbb{Z}$, and any permutation $\pi\in\Pi_{k}$. While second-order exchangeability, just as second-order stationarity, is mathematically sufficient for our needs, “full” exchangeability and “full” stationarity are the most natural and intuitively justifiable assumptions here. Denote by ${\mathcal{C}}_{*}\subset\mathcal{C}$ the class of second-order stationary second-order exchangeable processes $\mathbf{X}=\big{\{}X_{it}|i\in{\mathbb{N}},t\in{\mathbb{Z}}\big{\}}$.
@@
The distribution ${\mathbb{P}}$ here plays the role of a nuisance. Statistical practice in such cases consists in conducting inference on the realization observed in Step II conditional on the (unobserved) result of Step I (see, e.g., Chapter 10 of Lehmann and Romano (2024)), so that ${\mathbb{P}}$ needs no further description. Under such conditional approach, the random distribution ${\rm P}_{\bf X}$ of the process ${\bf X}\in{\mathcal{C}}_{*}$ of which the observed panel is a finite realization is treated as unknown but fixed. Hence ${\bm{\Sigma}}^{(n)}({\theta})$, ${\bm{\Gamma}}^{(n)}$, and their eigenvalues are random variables treated as unknown but fixed quantities on which conditions as (3.4) or (4.3) are imposed: under unconditional form, these conditions actually should be imposed $\mathbb{P}$-almost surely.

### v3

@@ 6 The Weak: De Mol, Giannone, and Reichlin, Onatski, Hallin and Liška, and Gersing, Rust, and Deistler
    Statically rate-weak factors in the static approximate factor model are introduced by De Mol et al. (2008) and Onatski (2012). They are related with the eigenvalues $\lambda^{(n)}_{\chi;j}$, $j=1,\ldots,r$, of the common component covariance matrix $\bm{\Gamma}_{\chi}^{(n)}$ which are diverging at sublinear rates, usually parametrized as $n^{\alpha_{j}}$ with $\alpha_{j}<1$. This is the most popular setting, see Lam and Yao (2012), Freyaldenhoven (2022), Uematsu and Yamagata (2022), and Bai and Ng (2023), among many others.
@@
    Dynamically rate-weak factors in the GDFM are considered by Barigozzi and Farnè (2024) who propose an estimator of a low-rank plus sparse spectral density matrix, under the assumption that the low-rank component has eigenvalues $\lambda^{(n)}_{\chi;j}(\theta)$, $j=1,\ldots,q$, diverging at sublinear rates $n^{\alpha_{j}}$ with $\alpha_{j}<1$, $\theta\text{-a.e. in }(-\pi,\pi]$.
@@
    Block-specific factors are introduced by Hallin and Liška (2011), under a GDFM approach, in the context of panels divided into subpanels or blocks. The terminology weakly common factors is used for dynamic factors that are dynamically common in some block(s) but dynamically idiosyncratic in some other(s). The associated spectral eigenvalues are assumed to be diverging with the dimension of the block but no rate is specified. Static approximate factor models with a hierarchical block structure are considered, for example, by Moench et al. (2013) but without any explicit mention of weak factors.
@@
    Finally, Gersing (2023) and Gersing et al. (2023) define weakly common components as the difference (at time $t$) between the dynamically common and the statically common components of $X_{it}$. Any (possibly infinite-dimensional) orthonormal basis of the space they are spanning (at time $t$) then can be considered a basis of weak factors at time $t$. This notion of weakness bears no relation to (a) and (b) above: to avoid confusion, call GRD-weak these weak factors.
@@ 6.1 Rate-weak factors: De Mol, Giannone and Reichlin, Onatski
As mentioned in the previous sections, as long as we have a widening, as $n\to\infty$, of the eigen-gap between the first $q$ or $r$ eigenvalues of the spectral density or covariance matrix of the observables $\mathbf{X}^{(n)}$ and the remaining $n-q$ or $n-r$ ones, the GDFM or the static approximate factor models are identified. Thus, for identification purposes, there is no need to assume any specific divergence rate for the divergent eigenvalues.
@@ 6.2 Weak factors in panels with block structure: Hallin and Liška
A similar approach could be taken, mutatis mutandis, with similar comments, in the context of static approximate factor models, yielding a refinement of the static factor model decomposition (2.1).
@@ 6.3 Weak factors are everywhere: Gersing, Rust, and Deistler
Under the intriguing title “Weak factors are everywhere,” Gersing et al. (2023), with the objective of reconciling the static and the dynamic approaches, show (their definitions slightly differ from the ones considered here) that the statically common space $\mathcal{H}_{\scriptscriptstyle\text{\rm stat com}}^{{\bf X}_{t}}$, for any $t$, is a subspace of the dynamically common one $\mathcal{H}_{\scriptscriptstyle\text{\rm dyn com}}^{{\bf X}}$ and call weakly common at time $t$ the elements of the difference $\mathcal{H}_{\scriptscriptstyle\text{\rm dyn com}}^{{\bf X}}\setminus\mathcal{H}_{\scriptscriptstyle\text{\rm stat com}}^{{\bf X}_{t}}$: we shall call them GRD-weak.
Let $\bf X$ be second-order stationary and satisfy both (3.4) and (5.3). This result by Gersing et al. (2023) naturally yields a new canonical factor decomposition, now involving three terms, of the form
@@
where $\chi_{it}^{\scriptscriptstyle\text{\rm weak}}$—the GRD-weakly common component of $X_{it}$—is the difference between the dynamically and statically common components $\chi_{it}^{\scriptscriptstyle\text{\rm dyn}}$ and $\chi_{it}^{\scriptscriptstyle\text{\rm stat}}$ (equivalently, the difference between the statically and dynamically idiosyncratic components $\xi_{it}^{\scriptscriptstyle\text{\rm stat}}$ and $\xi_{it}^{\scriptscriptstyle\text{\rm dyn}}$) of $X_{it}$. Denoting by ${\mathcal{H}}^{{\bf X}_{t}}_{\scriptscriptstyle\text{\rm weak}}$ the Hilbert space generated by $\mathcal{H}_{\scriptscriptstyle\text{\rm dyn com}}^{{\bf X}}\setminus\mathcal{H}_{\scriptscriptstyle\text{\rm stat com}}^{{\bf X}_{t}}$ (which depends on $t$) , we have, for all $t\in{\mathbb{Z}}$, the inclusions ${\mathcal{H}}^{{\bf X}_{t}}_{\scriptscriptstyle\text{\rm weak}}\subseteq\leavevmode\nobreak\ \!\mathcal{H}_{\scriptscriptstyle\text{\rm stat idio}}^{{\bf X}_{t}}$ and ${\mathcal{H}}^{{\bf X}_{t}}_{\scriptscriptstyle\text{\rm weak}}\subseteq\mathcal{H}_{\scriptscriptstyle\text{\rm dyn com}}^{{\bf X}}$, hence
$\mathcal{H}_{\scriptscriptstyle\text{\rm stat com}}^{{\bf X}_{t}}{\perp}_{t}\,\,{\mathcal{H}}^{{\bf X}_{t}}_{\scriptscriptstyle\text{\rm weak}}\perp\mathcal{H}_{\scriptscriptstyle\text{\rm dyn idio}}^{{\bf X}},\quad t\in{\mathbb{Z}}$ (6.2)
where $\perp$ and tsubscriptperpendicular-${\perp}_{t}$ mean orthogonal at all leads and lags and orthogonal at time $t$, respectively.
@@
This is a very ingenious finding, which shows how the GDFM (as defined in (3.5)) is overarching the static model (as defined in (2.1)) and, with the additional assumption (5.3), strictly nests the popular static approximate model of Chamberlain (1983) and Chamberlain and Rothschild (1983). The fact that the dynamically common space $\mathcal{H}_{\scriptscriptstyle\text{\rm dyn com}}^{{\bf X}}$ incorporates the GRD-weakly common components which, in the static approach, are idiosyncratic explains the empirical finding (Forni et al., 2018) that the GDFM performs better in terms, e.g., of forecasting, than the static model (2.1) of Chamberlain and Rothschild even when the assumptions of the latter are satisfied. See Section 9 for major further consequences of (6.1) and the orthogonality properties (6.2) on the relative forecasting performance of the static and the GDFM approaches.
Note that, from our definitions of the dynamically common and statically common-at-time-$t$spaces $\mathcal{H}_{\scriptscriptstyle\text{\rm dyn com}}^{{\bf X}}$ and $\mathcal{H}_{\scriptscriptstyle\text{\rm stat com}}^{{\bf X}_{t}}$ (which slightly differ from the Gersing et al. (2023) ones), the inclusion $\mathcal{H}_{\scriptscriptstyle\text{\rm stat com}}^{{\bf X}_{t}}\subseteq\mathcal{H}_{\scriptscriptstyle\text{\rm dyn com}}^{{\bf X}}$ trivially holds for any $t\in\mathbb{Z}$ so that, under the unique assumption of second-order stationarity, $X_{it}$ for all $i$ and all $t$ admits the canonical decomposition (6.1).
@@
In this canonical decomposition (6.1), $\chi_{it}^{\scriptscriptstyle\text{\rm weak}}$, being in the dynamically common space $\mathcal{H}_{\scriptscriptstyle\text{\rm dyn com}}^{{\bf X}}$, is orthogonal, at all leads and lags, to the dynamically idiosyncratic component ${\xi_{it}^{\scriptscriptstyle\text{\rm dyn}}}$; being in the statically idiosyncratic-at-time-$t$ space, $\chi_{it}^{\scriptscriptstyle\text{\rm weak}}$ is also orthogonal to the statically common-at-time-$t$ component ${\chi_{it}^{\scriptscriptstyle\text{\rm stat}}}$ (but not necessarily its leads and lags); that ${\chi_{it}^{\scriptscriptstyle\text{\rm stat}}}$ might even be a.s. zero (yielding a statically purely idiosyncratic process). Therefore, $\chi_{it}^{\scriptscriptstyle\text{\rm weak}}$ cannot be a static factor, neither rate-strong nor rate-weak, and typically consists of non-pervasive leading or lagging values of the (rate-weak or rate-strong) dynamic factors or combinations thereof. This concept of GRD-weakly common component, thus, is unrelated to the concepts of rate-weak factors discussed in Sections 6.1 and 6.2.
@@ 7 Exchangeability: weak factors are nowhere!
When no block structure is present in the data, one may wonder whether the concept of rate-weak factors is really relevant in practice.
An all too often overlooked feature of panel data is that, unlike time-ordering, their cross-sectional ordering is entirely arbitrary. Very often, some alphabetical ordering is adopted, a choice that should not have any impact on the analysis. An observed panel, actually, should be treated as the equivalence class of its $n!$ cross-sectional permutations and all statistical procedures, all forecasts, all inferential conclusions, should be invariant or equivariant under these cross-sectional permutations.
In this section, for simplicity of notation, we only consider the static case. The same arguments can be made verbatim for the dynamic case at the price of heavier notation.
@@ 7.1 Factors and cross-sectional permutations: intuition
The left-hand plot in Figure 1 shows the evolution, for increasing $n=1,\ldots,n_{0}$, of the first eigenvalue $\lambda_{X;1}^{(n)}$ (normalized by $\lambda_{X;1}^{(n_{0})}$) of the $(n\times n)$ lag-zero covariance matrix ${\bm{\Gamma}}^{(n)}_{X}$ in the case of three cross-sectional permutations of the $X_{it}$’s. Let $B_{(i)}$, $i=1,\ldots,n_{0}$ denote the $i$th order statistic of the loadings: $B_{(1)}\leq B_{(2)}\leq\leavevmode\nobreak\ \!\ldots\leavevmode\nobreak\ \!\leq\leavevmode\nobreak\ \!B_{(n_{0})}$. The first permutation (dashed line) is ordering the cross-sectional items by increasing values of the loadings ($B_{i}=B_{(i)}$); the second permutation (dashed-dotted line) orders them by decreasing values ($B_{i}=B_{(n_{0}-i+1)}$); the third one (solid line) is alternating small and large values of the same ($B_{1}=B_{(n_{0})},B_{2}=B_{(1)},B_{3}=B_{(n_{0}-1)},\ldots,B_{n_{0}}=B_{(n_{0}/2)}$ for even $n_{0}$). From a naïve point of view, one would be tempted by a superlinear extrapolation in the first case (hence, a rate-superstrong factor) and a sublinear one (a rate-weak factor) in the second case, whereas a linear extrapolation (a rate-strong factor) looks quite reasonable in the third case. The three cases, however, are dealing with the same panel.
@@ 7.2 Factors and cross-sectional permutations: exchangeability
Distributional symmetries and group invariance properties are among the most convincing and forceful arguments giving statistical inference problems a structure which, when it exists, dramatically impacts their solutions. Invariance arguments have a long history in mathematical statistics (see Chapter 6 in Lehmann and Romano (2024)); their pertinence, in a broad variety of problems, has been highlighted recently in Austern and Orbanz (2022).
Stationarity in time is one of those fundamental invariance properties and, under our fully nonparametric approach (no parameters), is the only assumption we are imposing in order to obtain a factor model decomposition. Namely, we assume that the observed panel is the finite $(n\times T)$ realization of a stochastic process $\bf X$ of the form $\big{\{}X_{it}|\,i\in{\mathbb{N}},t\in{\mathbb{Z}}\big{\}}$ in the class $\mathcal{C}$ of real-valued second-order time-stationary processes indexed on ${\mathbb{N}}\times{\mathbb{Z}}$. Second-order stationarity imposes a restriction (stationarity) on time dependence, namely,
@@
The stochastic process $\big{\{}X_{it}|i\in{\mathbb{N}},t\in{\mathbb{Z}}\big{\}}$ is (second-order) cross-sectionally exchangeable—that is, for any $k\in{\mathbb{N}}$, any $k$-tuple $i_{1},\ldots,i_{k}$, and any permutation $\pi$ in the set $\Pi_{k}$ of the permutations of the integers $(1,\ldots,k)$, the distributions (the second-order moments) of the $k$-dimensional stochastic subprocesses
$\big{\{}(X_{i_{1}t},\ldots,X_{i_{k}t})|\,t\in{\mathbb{Z}}\big{\}}\quad\text{and}\quad\big{\{}(X_{i_{{\pi(1)}}t},\ldots,X_{i_{{\pi(k)}}t})|\,t\in{\mathbb{Z}}\big{\}}$
@@
Clearly, exchangeability implies second-order exchangeability, viz. ${\rm E}[X_{i_{\ell}t}X_{i_{m}t^{\prime}}]={\rm E}[X_{i_{\pi(\ell)}t}X_{i_{\pi(m)}t^{\prime}}]$ for any $\ell$ and $m$ in $(1,\ldots,k)$, any $t$ and $t^{\prime}$ in $\mathbb{Z}$, and any permutation $\pi\in\Pi_{k}$. While second-order exchangeability, just as second-order stationarity, is mathematically sufficient for our needs, “full” exchangeability and “full” stationarity are the most natural and intuitively justifiable assumptions here. Denote by ${\mathcal{C}}_{*}\subset\mathcal{C}$ the class of second-order stationary second-order exchangeable processes $\mathbf{X}=\big{\{}X_{it}|\,i\in{\mathbb{N}},t\in{\mathbb{Z}}\big{\}}$.
@@
The distribution ${\mathbb{P}}$ here plays the role of a nuisance. Statistical practice in such cases consists in conducting inference on the realization observed in Step II conditional on the (unobserved) result of Step I (see, e.g., Chapter 10 of Lehmann and Romano (2024)), so that ${\mathbb{P}}$ needs no further description. Under such conditional approach, the random distribution ${\rm P}_{\bf X}$ of the process ${\bf X}\in{\mathcal{C}}_{*}$ of which the observed panel is a finite realization is treated as unknown but fixed. Hence ${\bm{\Sigma}}^{(n)}({\theta})$, ${\bm{\Gamma}}^{(n)}$, and their eigenvalues are random variables treated as unknown but fixed quantities on which conditions such as (3.4) or (5.3) are imposed: under unconditional form, these conditions actually should be imposed $\mathbb{P}$-almost surely.